A: A monopolist faces the following demand curve, marginal revenue curve, total cost curve for its product:

Q=3500-5p

MR= 250-Q

TC=15Q

MC=100

What level of output maximizes total revenue?

What is the profit-maximizing level of output?

What is the profit-maximizing price?

How much profit does the monopolist earn?

Suppose that a tax of $10 for each unit produced is imposed by the state government.

What is the profit-maximizing level of output?
Answer #1

Total revenue manim nising an autputi demand curne, 3500-5P ginen, g >> 5P = 3500 - 9 -> p = P = 700 - 0.2 from ☺ , tatal ren

prues profit maximising aufpect and at manimum profit, TORTE MR=me => 250 250-9 = >9* = 150 100 and profil manimining preie =auchen $10 ton is imposed per unit: The new met = old me + 10 100 + 10 Reme how profil maximising bul: MR = me* 250-9= 110 8